OpenStudy (anonymous):

2.)Suppose a population of 160 crickets doubles in size every month. The function f(x)=160*2^x 2,684,354,560 crickets 640 crickets 7.680 crickets 655,360 crickets 3.) Suppose an investment of $2,300 doubles in value every decade. The function f(x)=2,300*2^x gives the value of the investment after x decades. How much is the investment worth after 2 decades ? $2,355,200 $92,000 $46,000 $9,200

OpenStudy (anonymous):

2. 12 months in a year. 24 months in 2 years. 160 * 2^24 = 2,684,354,560 3. 2^2 = (2)(2) = 4 So $2300*4 = 9,200

OpenStudy (anonymous):

Thanks ^.^

OpenStudy (anonymous):

You're welcome!! Xo!

OpenStudy (anonymous):

You really helped a lot. (:

OpenStudy (anonymous):

Whenever you need help, just come by me and message me anytime!! Xo!

OpenStudy (anonymous):

OK thnx so much I appreciate it :)